OpenStudy (anonymous):

Find the slope of the line. y = 2/5 x + 8

OpenStudy (shamil98):

y = mx +b m = slope so your slope is.

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

-.4

OpenStudy (anonymous):

y= 2/5x+8 b= 8, so m=?

OpenStudy (anonymous):

-4 is not an answer

OpenStudy (anonymous):

.4 is the decimal form of 2/5 which is m

OpenStudy (anonymous):

a.2/5 b. -2/5 c. 8

OpenStudy (anonymous):

y=mx+b (m=slope, b=yintercept) y=2/5x+8 there is a pattern there. there fore the slope is 2/5. because it is m

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Slope = m, so the slope is 2/5 because mx+b makes m the slope, so 2/5 is the slope.

OpenStudy (anonymous):

.4 is equal to 2/5 also, so @stocks301 was not wrong.

OpenStudy (anonymous):

when you have an equation that is in the term of y=mx+b and you can recognize it more often that not the slope is right in front of the x (: